# Resummation-based Quantum Monte Carlo for Entanglement Entropy Computation

### v5

@@ Appendix A Performance of ResumEE
As mentioned in the main text, ResumEE method exponentially improves the sampling efficiency of the traditional entanglement entropy estimator $P_{\emptyset\to A}$ (and $P_{A\to\emptyset}$) through its ability to compress $N^{n_{l}}$ configurations into one compact, uncoloblack loop configuration, where $n_{l}$ denotes the number of loops in an SSE configuration. To be more specific, only those loops at the imaginary time edges contribute to the evaluation of $P$. Let $\bar{n}_{l}$ denote the number of loops at the imaginary time edge. It is shown explicitly in Fig. 6 that $\bar{n}_{l}(L,N)$ holds a power law relation with the linear system size $L$ (we keep $\beta=L$ to approach the ground state) and $N$, which demonstrates the exponential speedup of ResumEE.
@@ Appendix C Incremental approach of EE computation
Figure 8: Schematic of the incremental entanglement regions. $A_{k}$, $A_{k+1}$ and $A_{k+2}$ are the intermediate entanglement regions in which the ratio in Eq. (9) are computed independently. Since the incremental regions are not very different from each other, the incremental ratios computed are well-conditioned, compablack with direct computation of $\langle P_{\emptyset\rightarrow A}\rangle$ and $\langle P_{A\rightarrow\emptyset}\rangle$.

### v6

@@ Appendix A Performance of ResumEE
As mentioned in the main text, ResumEE method exponentially improves the sampling efficiency of the traditional entanglement entropy estimator $P_{\emptyset\to A}$ (and $P_{A\to\emptyset}$) through its ability to compress $N^{n_{l}}$ configurations into one compact, uncolored loop configuration, where $n_{l}$ denotes the number of loops in an SSE configuration. To be more specific, only those loops at the imaginary time edges contribute to the evaluation of $P$. Let $\bar{n}_{l}$ denote the number of loops at the imaginary time edge. It is shown explicitly in Fig. 6 that $\bar{n}_{l}(L,N)$ holds a power law relation with the linear system size $L$ (we keep $\beta=L$ to approach the ground state) and $N$, which demonstrates the exponential speedup of ResumEE.
@@ Appendix C Incremental approach of EE computation
Figure 8: Schematic of the incremental entanglement regions. $A_{k}$, $A_{k+1}$ and $A_{k+2}$ are the intermediate entanglement regions in which the ratio in Eq. (9) are computed independently. Since the incremental regions are not very different from each other, the incremental ratios computed are well-conditioned, compared with direct computation of $\langle P_{\emptyset\rightarrow A}\rangle$ and $\langle P_{A\rightarrow\emptyset}\rangle$.